OpenStudy (anonymous):

How would i check this in my calculator if it is correct or not. log(8) x + log(8) (x+30) = 2 log(8) (2x+30) = 2 8^2 = 2x+30 64 = 2x+30 34 = 2x 17 = x If i wanted to plug it back in.

OpenStudy (anonymous):

wrong!!... log add means times...

OpenStudy (anonymous):

log(8)x + log(8)(x+30)=2 log(8)[x(x+30)]=2 x(x+30)=8^2 x^2+30x=64 x^2+30x-64=0 (x+32)(x-2)=0 x+32=0 or x-2=0 x=-32 or x=2 (rejected) (in log x must be positive number...) so x=2

OpenStudy (anonymous):

sorry i hv a network problem

OpenStudy (anonymous):

u can try.. use your calculator and try 1. log(-any number)(+any number) 2. 1og(+any number)(-any number)

OpenStudy (anonymous):

u will see that cant use...

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

^^